OpenStudy (anonymous):

solve 15^-x=24. Round to the nearest hundredth

OpenStudy (anonymous):

x ≈ 2.45 x ≈ -1.29 x ≈ -1.17 x ≈ 1.67

OpenStudy (helder_edwin):

you have \[ \large 15^{-x}=24 \] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

log the both sides

OpenStudy (helder_edwin):

then take logarithms of base 15. so \[ \large \log_{15}15^{-x}=\log_{15}24 \] \[ \large -x=\log_{15}24 \]

OpenStudy (anonymous):

c?

OpenStudy (helder_edwin):

finally \[ \large \log_{15}24=\frac{\log 24}{\log 15} \] use a calculator for this

OpenStudy (helder_edwin):

you got it right!

OpenStudy (helder_edwin):

but use your calculator to be sure

OpenStudy (anonymous):

thanks :)